### v1

# MemEvo: Memory-Evolving Incremental Multi-view Clustering

Abstract
Incremental multi-view clustering aims to achieve stable clustering results while addressing the stability-plasticity dilemma (SPD) in incremental views. At the core of SPD is the challenge that the model must have enough plasticity to quickly adapt to new data, while maintaining sufficient stability to consolidate long-term knowledge and prevent catastrophic forgetting. Inspired by the hippocampal-prefrontal cortex collaborative memory mechanism in neuroscience, we propose a Memory-Evolving Incremental Multi-view Clustering method (MemEvo) to achieve this balance. First, we propose a hippocampus-inspired view alignment module that captures the gain information of new views by aligning structures in continuous representations. Second, we introduce a cognitive forgetting mechanism that simulates the decay patterns of human memory to modulate the weights of historical knowledge. Additionally, we design a prefrontal cortex-inspired knowledge consolidation memory module that leverages temporal tensor stability to gradually consolidate historical knowledge. By integrating these modules, MemEvo achieves strong knowledge retention capabilities in scenarios with a growing number of views. Extensive experiments demonstrate that MemEvo exhibits remarkable advantages over existing state-of-the-art methods.
Introduction
Recent advancements in multimedia technology have significantly transformed data analysis paradigms, where Multi-View Clustering (MVC) has demonstrated notable advantages in unsupervised scenarios by integrating heterogeneous data sources (Du et al. 2025; Wen et al. 2024; Fu et al. 2024). Traditional MVC methods have achieved significant success in processing static, complete datasets, as they require a complete set of views to achieve effective knowledge fusion (Du et al. 2024). However, data is not static but evolves dynamically over time in many real-world applications. For example, medical monitoring may extend from clinical measurements to imaging scan data, and multi-view data facilitates tracking of patient progress. This incremental property poses a major challenge to traditional methods, driving incremental multi-view clustering (IMVC) to become a key research direction. IMVC aims to efficiently update clustering results when new views arrive, thereby avoiding retraining the model from scratch. However, its core challenge is to resolve a fundamental contradiction: the Stability-Plasticity Dilemma (SPD) (Wu, Gong, and Li 2021). On the one hand, the model must possess sufficient plasticity to enable rapid learning and adaptation to new patterns and knowledge contained in new data. If the model is too rigid, it will be unable to capture dynamic changes in the data, leading to a rapid decline in performance. On the other hand, the model must also maintain sufficient stability to consolidate long-term accumulated knowledge and avoid catastrophic forgetting when learning new information. Therefore, an ideal IMVC model must achieve a balance between these two aspects.
Recent incremental multi-view clustering (IMVC) methods have shown significant progress. For example, Wan et al. (Wan et al. 2022) dynamically update a consensus partition matrix by transferring knowledge from historical views, enabling incremental learning. Yan et al. (Yan et al. 2024) further introduced a category memory library to store learned categories, while Qu et al. (Qu et al. 2024) proposed a lightweight framework that retains and jointly optimizes the consensus anchor graph of previous views. Despite existing IMVC methods having made valuable explorations, existing IMVC methods often address only specific aspects of SPD. Some prioritize plasticity, rapidly integrating new data at the risk of catastrophic forgetting, while others emphasize stability through regularization, potentially hindering adaptation to new views. A key limitation is that existing methods fail to balance old knowledge retention and new knowledge integration, which causes performance bottlenecks as views continuously grow. To address this dilemma, we drew inspiration from the memory system of the human brain. Cognitive science research has found that the human brain, particularly the collaborative memory system between the hippocampus and the prefrontal cortex, serves as a paradigm for balancing plasticity and stability. The hippocampus can quickly encode new episodic memories and is highly plastic (Arani, Sarfraz, and Zonooz 2022). In contrast, the prefrontal cortex integrates and consolidates these temporary memories into long-term knowledge (highly stable) through a slow, gradual process. This complementary mechanism of rapid learning and slow consolidation enables humans to continue learning without easily forgetting the past. Furthermore, the acquisition of knowledge is not merely an accumulation of information. The brain also employs a complex active adaptive forgetting mechanism to discard outdated information and maintain cognitive efficiency (Anderson and Hulbert 2021). This forgetting is not a passive process of memory decay but an active process of neural remodeling that occurs immediately after learning behavior is completed (Hardt, Nader, and Nadel 2013).
Inspired by this, we propose a Memory-Evolving Incremental Multi-view Clustering method (MemEvo) that aims to solve the plasticity-stability dilemma. We should clarify that MemEvo is not designed to create a biologically accurate brain simulation. Instead, we borrow from their division of roles between plastic learning and stable integration to build a more effective incremental clustering algorithm. To this end, we have designed a collaborative framework comprising three core modules. First, we propose a hippocampus-inspired view alignment module, which effectively captures incremental information from new views by aligning structures in the continuous representation space, thereby ensuring plasticity of our model. Second, we introduce a cognitive forgetting mechanism. This mechanism simulates the decay patterns of human memory, selectively forgetting historical information by adjusting weights. Finally, we designed a prefrontal cortex-inspired knowledge consolidation memory module, which leverages the stability of temporal tensors to gradually consolidate historical knowledge, ensuring the stability of the model. By combining these three core modules, MemEvo achieves remarkable knowledge retention capabilities in view-incremental scenes. Our main contributions are summarized as follows:
  • A Novel Neuroscience-Inspired Framework: This paper first introduces the hippocampus-prefrontal cortex collaborative memory model into incremental multi-view clustering, providing a new perspective for solving the stability-plasticity dilemma.
  • Collaborative Module Design: We designed a collaborative optimization framework that organically integrates three core modules: view alignment, cognitive forgetting, and knowledge consolidation memory, achieving efficient and stable incremental learning.
  • State-of-the-Art Performance: Extensive experiments show that our proposed method outperforms existing incremental clustering methods.
Related Work
Incremental multi-view clustering
Incremental Multi-View Clustering (IMVC) continuously integrates new views while retaining historical knowledge, thereby achieving stable clustering. Existing literature achieves incremental updates by maintaining historical knowledge carriers (Wan et al. 2022; Lu et al. 2024; Wan et al. 2024; Chen et al. 2025). For example, CMVC (Wan et al. 2022) and CAC (Yan et al. 2024) maintain consensus partition matrices to achieve knowledge reuse. FDMVC (Lu et al. 2024) maintains semantically consistent representations for updates. LAIMVC (Qu et al. 2024) maintains consensus anchor graphs and aligns new view anchors through permutation matrices. Consistent with existing research, MemEvo is based on the following assumptions: each view contains complete class information, and class growth scenarios are not considered. Based on this, we focus on how to simulate cognitive mechanisms to achieve efficient capture of multi-view information.
Nonconvex Tensor rank substitution
Non-convex methods have been proven (Zhong et al. 2015) to approximate the tensor rank more accurately, and thus have become the current dominant technical paradigm in traditional multi-view clustering (Guo et al. 2022; Long et al. 2025; Ji and Feng 2025). Let $\{\mathbf{X}_{t}\}_{t=1}^{T}\in\mathbb{R}^{n\times d_{t}}$ be multi-view data with $T$ views, where $d_{t}$ is the dimension of the view and $n$ is the sample number. We can optimize the following objective function to achieve our goal.
$\min_{\mathbf{A}_{t},\mathbf{Z}_{t}}\sum_{t=1}^{T}\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}\|_{2,1}+\lambda\|\mathcal{Z}\|_{low},s.t.\mathcal{Z}=\Phi(\mathbf{Z}_{1},..,\mathbf{Z}_{t})$ (1)
where $\mathbf{A}_{t}\in\mathbb{R}^{m\times d_{t}}$ (where $m$ is the latent space dimension) is the basic matrix and $\mathbf{Z}_{t}$ is the representation for $t$-view. Besides, $\Phi(\cdot)$ constructs the tensor $\mathcal{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ by stacking representation $\mathbf{Z}_{t}\in\mathbb{R}^{n\times m}$ from each view into a three-order tensor, $\|\cdot\|_{low}$ is the rank constraint term. Here, a commonly used nonconvex substitution function (Lu et al. 2015) is introduced, which is defined as:
Definition 1.
For tensor $\mathcal{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, the Alternative Rank Minimization Regularization (ARMR) is defined as:
$\|\mathcal{Z}\|_{ARMR}=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}\sum_{i=1}^{min(n_{1},n_{2})}\left(\frac{1-e^{-\mathcal{S}_{f}^{k}(i,i)}}{1+e^{-\mathcal{S}_{f}^{k}(i,i)}}\right)$ (2)
With the Fast Fourier Transform (FFT) along the third dimension, we can obtain $\mathcal{Z}_{f}^{k}$ by $fft(\mathcal{Z})=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}\mathcal{Z}_{f}^{k}$. Then, we perform tensor singular value decomposition(t-SVD) to obtain $\mathcal{Z}_{f}^{k}=\mathcal{U}*\mathcal{S}_{f}^{k}*\mathcal{V}^{T}$ (Kilmer et al. 2013).
Figure 1: The framework of MemEvo. The model achieves the retention of historical knowledge and the learning of new knowledge through three core modules: cognitive forgetting, view alignment, and knowledge consolidation.
The Proposed method
Overview and Motivation
In incremental multi-view clustering, maintaining a balance between learning new knowledge (plasticity) and retaining existing knowledge (stability) is a major challenge. To address this, we draw inspiration from the human brain’s memory system. Specifically, we have simulated the collaborative memory mechanism between the hippocampus and the prefrontal cortex. As shown in Figure 1, MemEvo contains three core components organized to address the Stability-Plasticity Dilemma: a hippocampus-inspired view alignment module, a cognitive forgetting mechanism, and a prefrontal cortex-inspired knowledge consolidation memory module.
Plasticity: New Knowledge Encoding & Alignment
When new view data arrives, the model must quickly learn and adapt to the new knowledge it contains. We simulate the rapid learning ability of the brain through incremental representation learning and view alignment.
Foundation: Incremental Representation Learning. First, we designed an incremental representation learning strategy to establish a foundation for subsequent memory operations. When a new data view $\mathbf{X}_{t}$ is received, we reconstruct the original data by a representation $\mathbf{Z}_{t}$ and an orthogonal basis matrix $\mathbf{A}_{t}$. This process can be expressed as reconstruction loss:
$\mathcal{L}_{recon}=\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}\|_{2,1},\quad s.t.\mathbf{A}_{t}\mathbf{A}_{t}^{T}=\mathbf{I}.$ (3)
where $\mathbf{A}_{t}\mathbf{A}_{t}^{T}=\mathbf{I}$ ensures that the basis vectors are uncorrelated (Chen et al. 2022). $\|\cdot\|_{2,1}$ is the $\ell_{2,1}$ norm, which is used to enhance the robustness of model (Nie et al. 2010).
View Alignment Module (VAM): Simulating the rapid association of the hippocampus. In cognitive science, the hippocampus serves as a temporary storage center for short-term memory, playing a key role in the initial encoding and integration of new and old information (McClelland, McNaughton, and Lampinen 2020). To simulate this rapid initial association, we propose a hippocampus-inspired view alignment module (VAM). This module aims to achieve initial interaction between new and old knowledge, enabling the model to adapt to past knowledge while learning new tasks. Specifically, we introduce an orthogonal matrix $\mathbf{P}_{t}\in\mathbb{R}^{m\times m}$ to align the current representation $\mathbf{Z}_{t}$ with the previous representation $\mathbf{Z}_{t-1}$. This process can be expressed as
$\mathcal{L}_{align}=\|\mathbf{Z}_{t}-{\mathbf{Z}_{t-1}\mathbf{P}_{t}}\|_{F}^{2},s.t.\mathbf{P}_{t}\mathbf{P}_{t}^{T}=\mathbf{I}.$ (4)
Although VAM uses a technically mature orthogonal alignment method, this choice is a principled abstraction of hippocampal. The hippocampus is characterized by its ability to rapidly encode new events with high plasticity. Our VAM module aligns the current view with the previous view, enabling rapid processing, which corresponds to the encoding function of the hippocampus.
Stability: Knowledge Evolution and Consolidation
Although plasticity enables systems to adapt to new information, it must be balanced with stability to prevent catastrophic forgetting of accumulated knowledge. To achieve this, we simulate knowledge evolution through cognitive forgetting and knowledge consolidation memory.
Cognitive Forgetting Module (CFM): Improving the Expression of Historical Knowledge. In the human brain, knowledge acquisition is not a simple linear accumulation, but rather involves the active forgetting mechanism constantly clearing out outdated information to maintain the efficient operation of the cognitive system. Based on Ebbinghaus’s forgetting theory (Ebbinghaus 1985), we have introduced a cognitive forgetting module to process view information from past moments. This theory states that memory decay follows a pattern of rapid decline initially, followed by a slower rate, so the contribution of historical knowledge should not be treated equally but should gradually diminish over time. Since the mathematical properties of power-law functions exhibit a trend of rapid decline followed by gradual reduction (Heathcote, Brown, and Mewhort 2000), we utilize this tool to model the decay process. Specifically, for past representation $\mathbf{Z}_{i}$ (where $i\in\{1,...,t-1\}$), we calculate weight $w_{i}^{(t)}$ based on its temporal relevance to the current step $t$ by power-law functions:
$w_{i}^{(t)}=\frac{(t-i)^{-\lambda}}{\sum_{j=1}^{t-1}(t-j)^{-\lambda}},\quad i\in\{1,\dots,t-1\}.$ (5)
where $\lambda>0$ controls the forgetting rate. For view $t$, the integration of historical knowledge can be written as:
$\mathbf{Z}_{hist}=\sum_{i=1}^{t-1}w_{i}^{(t)}\mathbf{Z}_{i}.$ (6)
CFM assigns higher weights to recent representations and allows the contributions of earlier representations to decay. This scheme simulates the pattern of memory evolution in the brain to improve the expression of historical knowledge.
Knowledge Consolidation Memory Module (KCM): Integration for Stable Long-Term Memory. To address the critical challenge of stability in continuous learning, we draw inspiration from the memory mechanism of the prefrontal cortex. In the human brain, the prefrontal cortex collaborates with the hippocampus to convert short-term knowledge into stable long-term memory. To simulate this collaborative mechanism, we propose a knowledge consolidation memory module (KCM). Specifically, we construct a temporal tensor $\mathcal{Z}\in\mathbb{R}^{n\times m\times 2}$ by stacking the historical representation $\mathbf{Z}_{hist}$ and the current representation $\mathbf{Z}_{t}$, thereby achieving high-order interaction between old and new knowledge. This process is expressed as
$\mathcal{Z}=stack(\mathbf{Z}_{hist},\mathbf{Z}_{t},3).$ (7)
where $\text{stack}(\cdot,\cdot,3)$ denotes stacking along the third dimension (where the third dimension has only two slices, among which $\mathcal{Z}(:,:,1)=\mathbf{Z}_{hist},\mathcal{Z}(:,:,2)=\mathbf{Z}_{t},$). Next, to enhance consistency and consolidate the merged knowledge into a low-rank structure, we applied alternative rank minimization regularization (ARMR), which is defined in Definition 2. The process is expressed as
$\mathcal{L}_{consolidate}=\|\mathcal{Z}\|_{ARMR},\ s.t.\mathcal{Z}=stack(\mathbf{Z}_{hist},\mathbf{Z}_{t},3).$ (8)
KCM combines historical and current representations into a higher-order tensor to simulate the integration of the prefrontal cortex. Meanwhile, low-rank approximation captures the most critical structural information in the data, aligning with the goal of memory consolidation. This process correlates with the properties of the prefrontal cortex, ultimately forming a stable representation of long-term knowledge.
The Overall Objective Function
By combining the above modules responsible for plasticity and stability, the optimization objective of MemEvo can be expressed as:
$\mathcal{L}_{MemEvo}=\mathcal{L}_{recon}+\alpha\mathcal{L}_{align}+\beta\mathcal{L}_{consolidate}.$ (9)
The final optimization problem is then formulated as
$\min_{\begin{subarray}{c}\mathbf{Z}_{t},\mathbf{A}_{t},\\ \mathbf{P}_{t},\mathcal{Z}\end{subarray}}$ $\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}\|_{2,1}+\alpha\|\mathbf{Z}_{t}-\mathbf{Z}_{t-1}\mathbf{P}_{t}\|_{F}^{2}+\beta\|\mathcal{Z}\|_{ARMR},$ (10)
$s.t.$ $\mathbf{Z}_{hist}=\sum_{i=1}^{t-1}\frac{(t-i)^{-\lambda}}{\sum_{j=1}^{t-1}(t-j)^{-\lambda}}\mathbf{Z}_{i},\mathbf{A}_{t}\mathbf{A}_{t}^{T}=\mathbf{I},$
$\mathcal{Z}=stack(\mathbf{Z}_{hist},\mathbf{Z}_{t},3),\mathbf{P}_{t}\mathbf{P}_{t}^{T}=\mathbf{I}.$
Model optimization
We solve Eq.(10) using the Alternating Direction Method of Multipliers (ADMM) (Boyd et al. 2011). It decomposes the original problem into easily solvable subproblems by introducing auxiliary variables and Lagrange multipliers. To facilitate the solution, we introduce auxiliary variables $\mathbf{E}_{t}$ and $\mathcal{M}$ instead of $\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}$ and $\mathcal{Z}$, respectively. Then, the Lagrangian function can be written as
$\min_{\mathbf{Z}_{t},\mathbf{A}_{t},\mathbf{P}_{t},\mathcal{Z}}\|\mathbf{E}_{t}\|_{2,1}+\alpha\|\mathbf{Z}_{t}-\mathbf{Z}_{t-1}\mathbf{P}_{t}\|_{F}^{2}+\beta\|\mathcal{M}\|_{ARMR}$ (11)
$+\frac{\mu}{2}\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}-\mathbf{E}_{t}+\frac{\mathbf{Y}_{t}}{\mu}\|_{F}^{2}+\frac{\rho}{2}\|\mathcal{Z}-\mathcal{M}+\frac{\mathcal{J}}{\rho}\|_{F}^{2}.$
where $\mathbf{Y}$ and $\mathcal{J}$ denotes the Lagrange multiplier, $\mu>0,\rho>0$ is the penalty parameter.
$\triangleright$ Initial view processing ($t=1$).
For the first view, no history information exists, so the objective function simplifies to the base view representation learning part:
$\min_{\mathbf{A}_{1},\mathbf{E}_{1},\mathbf{Z}_{1}}\frac{\mu}{2}\|\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}-\mathbf{E}_{1}+\frac{\mathbf{Y}_{1}}{\mu}\|_{F}^{2}+\|\mathbf{E}_{1}\|_{2,1},$ (12)
$s.t.\mathbf{A}_{1}\mathbf{A}_{1}^{T}=\mathbf{I}.$
Update $\mathbf{A}_{1}$.
Fixing $\mathbf{Z}_{1}$, $\mathbf{E}_{1}$, we can obtain $\mathbf{A}_{1}$ by:
$\mathop{\arg\min}\limits_{\mathbf{A}_{1}}\frac{\mu}{2}\|\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}-\mathbf{E}_{1}+\frac{\mathbf{Y}_{1}}{\mu}\|_{F}^{2}\ s.t.\mathbf{A}_{1}\mathbf{A}_{1}^{T}=\mathbf{I}_{m}.$ (13)
This is a standard Orthogonal Procrustes Problem (Wen et al. 2018). Let $\mathbf{B}=\mu\mathbf{X}_{1}-\mu\mathbf{E}_{1}+\mathbf{Y}_{1}$, by performing a singular value decomposition (SVD) of $\mathbf{B}$ to obtain $\mathbf{B}=\mathbf{U}_{A}\mathbf{S}\mathbf{V}_{A}^{T}$, then the closed solution of $\mathbf{A}_{1}$ is $\mathbf{A}_{1}=\mathbf{U}_{A}\mathbf{V}_{A}^{T}$. $\mathbf{U}_{A}$ and $\mathbf{V}_{A}$ are the left and right singular value matrices of $\mathbf{B}\mathbf{Z}_{1}^{T}$.
Update $\mathbf{Z}_{1}$.
Fixing the other variables, the subproblem of $\mathbf{Z}_{1}$ can be solved in the following way
$\mathop{\arg\min}\limits_{\mathbf{Z}_{1}}\frac{\mu}{2}\|\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}-\mathbf{E}_{1}+\frac{\mathbf{Y}_{1}}{\mu}\|_{F}^{2}.$ (14)
The problem can be solved in closed form by taking the partial derivatives, and the solution is
$\mathbf{Z}_{1}=(\mathbf{X}_{1}-\mathbf{E}_{1}+\frac{\mathbf{Y}_{1}}{\mu})\mathbf{A}_{1}^{T}.$ (15)
Update $\mathbf{E}_{1}$.
Fixing other variables, $\mathbf{E}_{1}$ can be obtained by
$\mathop{\arg\min}\limits_{\mathbf{E}_{1}}\frac{\mu}{2}\|\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}-\mathbf{E}_{1}+\frac{\mathbf{Y}_{1}}{\mu}\|_{F}^{2}+\|\mathbf{E}_{1}\|_{2,1}.$ (16)
The solution of Eq.(LABEL:E) can be obtained by minimizing the thresholding operator in (Zhang et al. 2015), which is
${(\mathbf{E}_{1})}_{(:,i)}=\left\{\begin{matrix}\frac{\|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}-\frac{1}{\mu}}{\|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}}{(\mathbf{C}_{1})}_{(:,i)}&if\ \|{(\mathbf{C}_{1})}_{(:,i)}\|_{2}>\frac{1}{\mu};\\ 0&Otherwise.\end{matrix}\right$. (17)
where $\mathbf{C}_{1}=\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}+\frac{\mathbf{Y}_{1}}{\mu}$
Update $\mathbf{Y}_{1}$ and $\mu$.
The Lagrange multipliers and penalty parameters can be updated by the following rules
$\mathbf{Y}_{1}=\mathbf{Y}_{1}+\mu(\mathbf{X}_{1}-\mathbf{Z}_{1}\mathbf{A}_{1}-\mathbf{E}_{1}),\mu=min(\delta\mu,\mu_{max})$ (18)
where $\delta>1$ to accelerate convergence, $\mu_{max}$ are the positive constants. $\mathbf{Z}_{1}$ will serve as the starting point and first historical memory for the subsequent optimization process.
$\triangleright$ Updates to subsequent views ($t>1$).
Update $\mathbf{A}_{t}$ and $\mathbf{E}_{t}$.
Similarly, the update rules for $\mathbf{A}_{t}$ and $\mathbf{E}_{t}$ are similar to the first view from Eq.(13) and Eq.(17). The solution is obtained by processing $t$ = 2 to $T$ in turn.
Update $\mathbf{P}_{t}$.
This subproblem can be written as
$\min_{\mathbf{P}_{t}}\alpha\|\mathbf{Z}_{t}-\mathbf{Z}_{t-1}\mathbf{P}_{t}\|_{F}^{2},s.t.\mathbf{P}_{t}\mathbf{P}_{t}^{T}=\mathbf{I}.$ (19)
Similar to solving for $\mathbf{A}_{t}$, we can perform SVD on $\mathbf{C}_{t}=\mathbf{Z}_{t}^{T}\mathbf{Z}_{t-1}$ to obtain $\mathbf{P}_{t}=\mathbf{U}_{p}\mathbf{V}_{p}^{T}$.
Update $\mathcal{M}$.
When $\mathbf{A}_{t},\mathbf{Z}_{t},\mathbf{P}_{t},\mathbf{E}_{t}$ are fixed, we can update $\mathcal{M}$ as follow:
$\min_{\mathcal{M}}\beta\|\mathcal{M}\|_{ARMR}+\frac{\rho}{2}\|\mathcal{M}-(\mathcal{Z}+\frac{\mathcal{J}}{\rho})\|_{F}^{2}$ (20)
This is a classic Convex-ConCave Procedure Problem (CCCP) (Yuille and Rangarajan 2001), and we can solve it by the following Lemma:
Lemma 1.
Let tensor $\mathcal{M}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ with t-SVD. The non-concave tensor rank minimization regularization meets:
$\mathop{\arg\min}\limits_{\mathcal{M}}\frac{\beta}{\rho}\|\mathcal{M}\|_{ARMR}+\frac{1}{2}\|\mathcal{X}-\mathcal{M}\|_{F}^{2}$ (21)
The optimal solution of Eq.(21) can be written as
$\tilde{\mathcal{X}}=\Gamma_{\frac{1}{\rho}}(\mathcal{M})=\mathcal{U}*ifft(f(\bar{\mathcal{S}_{f}^{k}})(i,i))*\mathcal{V}^{T}.$ (22)
where ifft denotes the Inverse Fast Fourier Transform. The function $f(\bar{\mathcal{S}_{f}^{k}})$ is a f-diagonal tensor, which is
$f(\bar{\mathcal{S}_{f}^{k}}(i,i),x)=\mathop{\arg\min}\limits_{x\geq 0}\frac{1}{2}(x-\mathcal{S}_{f}^{k}(i,i))^{2}+\left(\frac{1-e^{-x}}{1+e^{-x}}\right)$ (23)
Eq. (23) is a CCCP problem, so we can use the difference of convex programming (Tao and An 1997) to obtain a closed-form solution as:
$f(\mathcal{S}_{f}^{k}(i,i))^{iter+1}=(\mathcal{S}_{f}^{k}(i,i)-\frac{\nabla f(\mathcal{S}_{f}^{k}(i,i))^{iter}}{\rho})_{+}$ (24)
where $(\cdot)_{+}$ denotes $max(\cdot,0)$, $\nabla f(\cdot)$ is the gradient of $f(\cdot)$.
Update $\mathbf{Z}_{t}$.
Fixing the other variables, the optimization function for $\mathbf{Z}_{t}$ can be written as
$\mathop{\arg\min}\limits_{\mathbf{Z}_{t}}\alpha\|\mathbf{Z}_{t}-\mathbf{Z}_{t-1}\mathbf{P}_{t}\|_{F}^{2}+\frac{\mu}{2}\|\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}-\mathbf{E}_{t}+\frac{\mathbf{Y}_{t}}{\mu}\|_{F}^{2}$ (25)
$+\frac{\rho}{2}\|\mathbf{Z}_{t}-\mathbf{M}_{t}+\frac{\mathbf{J}_{t}}{\rho}\|_{F}^{2}.$
This is a convex quadratic function whose optimal solution can be obtained by setting its derivative for $\mathbf{Z}_{t}$ to zero. Its closed solution can be written as
$\mathbf{Z}_{t}$ $=((\mu+2\alpha+\rho)\mathbf{I})^{-1}(\mu(\mathbf{X}_{t}-\mathbf{E}_{t}+\frac{\mathbf{Y}_{t}}{\mu})\mathbf{A}_{t}^{T}$ (26)
$+2\alpha\mathbf{Z}_{t-1}\mathbf{P}_{t}+\rho\mathbf{M}_{t}-\mathbf{J}_{t}).$
Updating other variables.
The Lagrange multipliers and penalty parameters can be updated by
$\mathbf{Y}_{t}=\mathbf{Y}_{t}+\mu(\mathbf{X}_{t}-\mathbf{Z}_{t}\mathbf{A}_{t}-\mathbf{E}_{t}),\mathcal{J}=\mathcal{J}+\rho(\mathcal{Z}-\mathcal{M}).$ (27)
$\mu=min(\delta\mu,\mu_{max}),\rho=min(\delta\rho,\rho_{max})$ (28)
where $\rho_{max}$ are the positive constants. The solution of our method is summarized in Algorithm 1, and the convergence analysis is shown in the supplementary material.
Algorithm 1 Updates to subsequent views ($t>1$).
Input: Incremental data $\{\mathbf{X}_{t}\}_{t=1}^{T}$, the latent dimension of basic matrix $m$, cluster number $k$.
Parameters: $\alpha,\beta$, $\lambda$
Output: After obtaining $\mathbf{Z}_{T}$ of the last view, we use $k$-means to obtain the clustering results.
1:Initialize: $\mathbf{A}_{t}$, $\mathbf{P}_{t}$, $\mathbf{E}_{t}$, $\mathbf{Z}_{t}$, $\mathbf{Y}_{t}$ = 0. $\mathcal{Z}$, $\mathcal{J}$, $\mathcal{M}$ = 0, $\mu$ = $\rho=10^{-4}$, $\mu_{max}$ = $\rho_{max}=10^{10}$, $\delta$ = 2.
2:while not converged do
3:  Update $\mathbf{Z}_{hist}$ and $\mathbf{A}_{t}$ by Eq.(6) and Eq. (13).
4:  Update $\mathbf{E}_{t}$ and $\mathbf{P}_{t}$ by Eq. (17) and Eq. (19).
5:  Update $\mathbf{Z}_{t}$ and $\mathcal{M}$ by Eq. (26) and Eq. (24).
6:  Update $\mathbf{Y}_{t}$ and $\mathcal{J}$ by Eq. (27).
7:  Update $\mu$ and $\rho$ by Eq. (28).
8:end while
Complexity Analysis
The computational complexity of our model is analyzed in two stages: the initial view processing ($t=1$) and the subsequent incremental updates ($t>1$). For the initial view, the primary computational costs arise from iteratively updating $\mathbf{A}_{1}$, $\mathbf{Z}_{1}$, and $\mathbf{E}_{1}$. The complexities for these updates are $\mathcal{O}(3dmn+dm^{2})$$\mathcal{O}(dmn)$$\mathcal{O}(dmn)$, respectively. Therefore, the complexity for the initial stage is dominated by $\mathcal{O}(dmn+dm^{2})$ per iteration. For each subsequent incremental step, the complexities for updating $\mathbf{Z}_{hist}$, $\mathbf{P}_{t}$, $\mathcal{M}$, $\mathbf{A}_{t}$, $\mathbf{E}_{t}$, and $\mathbf{Z}_{t}$ are $\mathcal{O}(tdn)$, $\mathcal{O}(m^{2}n+m^{3})$, $\mathcal{O}(m^{3}n+2mntlogt)$, $\mathcal{O}(dmn+dm^{2})$, $\mathcal{O}(dmn)$ and $\mathcal{O}(dmn+m^{2}n)$, respectively. Consequently, the overall complexity per iteration for an incremental step is $\mathcal{O}(tdn+3dmn+m^{3}n+dm^{2}+2m^{2}n+m^{3}+2mntlogt)$. Since $d\ll n$ and $m\ll n$, the computational cost grows approximately linearly with $n$. This linear scalability demonstrates the suitability of our model for processing large-scale datasets.
Experiments
Datasets and Evaluation. To assess the efficacy of the proposed method, we conducted experiments on six benchmark multi-view datasets: ProteinFold, Flower17, GRAZ02, handwritten, Caltech101-all, and MNIST. These datasets exhibit considerable variation in scale, with sample sizes ranging from 694 to 50,000 instances, and the maximum number of views is 27. Besides, three clustering metrics were used for performance evaluation: clustering accuracy (ACC), normalized mutual information (NMI), and adjusted rand index (ARI). Higher value indicates better performance.
Method ProteinFold Flower17 GRAZ02 handwritten Caltech101-all MNIST
ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI
Non-Incremental Methods
FMVACC 27.38 36.50 10.64 49.64 50.27 31.89 52.95 17.55 16.70 91.36 87.65 84.82 23.03 38.93 15.81 53.83 48.98 38.40
AWMVC 34.29 42.93 16.05 56.25 55.34 38.84 37.33 5.91 4.95 74.95 71.77 62.91 27.35 49.65 23.27 54.62 47.03 38.15
3AMVC 29.26 37.33 11.78 45.12 43.30 24.58 39.17 6.01 5.62 90.15 87.05 83.92 22.89 46.05 17.95 38.78 38.58 20.65
NpGC 37.18 43.66 17.32 58.01 57.44 40.34 42.68 8.64 8.17 89.35 80.71 78.10 28.37 49.60 24.17 55.30 47.79 38.77
CAMVC 33.14 42.80 16.67 53.01 50.12 32.77 43.22 13.08 10.22 91.95 84.80 83.11 26.41 47.74 22.18 39.71 36.30 23.34
IWTSN 44.09 70.59 36.27 93.27 92.93 88.79 70.08 34.51 32.91 99.65 99.10 99.22 60.94 87.55 43.85 99.63 98.97 99.18
DSCMC 29.16 39.02 12.80 34.86 44.89 23.59 38.73 5.64 5.17 71.90 68.73 58.76 24.05 42.93 18.00 67.22 59.24 52.35
LMTC 31.63 39.50 14.29 61.38 60.97 44.09 42.35 7.71 7.36 95.10 89.23 89.40 25.20 46.53 39.08 60.34 59.51 48.37
Incremental Methods
CMVC 31.76 40.55 14.02 55.61 55.02 38.27 48.60 12.22 12.01 89.85 80.25 78.96 24.46 46.88 20.54 60.77 53.30 44.55
LAIMVC 31.89 41.32 15.27 40.37 43.83 25.56 40.88 6.95 6.42 66.98 56.46 48.40 20.34 43.40 16.42 54.37 50.40 39.78
FDMVC 36.51 47.27 19.21 47.17 50.43 31.02 43.78 9.38 8.85 79.68 77.31 69.65 21.98 44.43 18.73 76.79 72.35 66.19
FCMVC 34.01 41.96 16.12 47.22 49.73 31.47 40.40 6.34 5.93 59.59 54.57 41.54 27.05 46.87 21.62 OM OM OM
Ours 67.15 82.02 58.15 94.85 94.27 90.56 92.28 80.45 80.52 99.95 99.86 99.89 67.10 89.16 56.97 99.87 99.60 99.73
Table 1: Performance results for six datasets compared to twelve state-of-the-art methods (”OM” stands for Out of Memory).
Comparative Methods. To demonstrate the effectiveness of our methods, we select the following methods for comparison. Traditional Methods: FMVACC (NeurIPS’22) (Wang et al. 2022), AWMVC (AAAI’23) (Wan et al. 2023), 3AMVC(MM’24)(Ma et al. 2024), NpGC (AAAI’24) (Yu et al. 2024), IWTSN (MM’24) (Sun et al. 2024), CAMVC (AAAI’24) (Zhang et al. 2024), DSCMC (PR’25) (Kong et al. 2025), LMTC (CVPR’25) (Liu et al. 2025). Incremental Methods: CMVC (MM’22) (Wan et al. 2022), LAIMVC (MM’24) (Qu et al. 2024), FDMVC (KBS’24) (Lu et al. 2024), FCMVC (TIP’24) (Wan et al. 2024).
Experiment Setup. All experiments were conducted on a personal computer with an Intel Core i9-13900K CPU and 64 GB of RAM. For all comparison methods, we adopted the parameter settings recommended in their original publications. Considering that $k$-means is more sensitive to initialization, we conduct 10 independent experiments and report the average of the clustering metrics to ensure fairness.
Experimental Results and Analysis
Table 1 shows the performance of the proposed method on six datasets with twelve comparison methods. 1) Despite being designed for incremental learning, the present method outperforms most traditional non-incremental methods. Notably, the proposed approach yields performance gains of 23.06% and 22.2% in ACC over the second-best method on ProteinFold and GRAZ02 datasets, respectively. 2) Our method demonstrates improvements over existing incremental methods (CMVC, LAIMVC, FDMVC, FCMVC). These improvements can be attributed to the balance of VAM and KCM, which addresses the plasticity-stability dilemma in incremental learning, thereby enabling efficient knowledge accumulation in incremental tasks. To enhance the validity of the results, we have provided code in the supplementary materials to facilitate reproduction.
Parametric Analysis
Sensitivity analysis of $\alpha$ and $\beta$.
To analyze the sensitivity of parameters $\alpha$ and $\beta$, we utilize the GRAZ02 and handwritten datasets as examples. With other parameters fixed, we investigate how variations within the interval $\{10^{-3},10^{-2},10^{-1},10^{0},10^{1}\}$ affect the accuracy (ACC) of our method. As shown in Figure 3(a), our method demonstrates low sensitivity to parameter variations when $\alpha$ and $\beta$ are within $\alpha,\beta\in\{10^{-3},...,10^{-1}\}$.
(a) Sensitivity analysis of $\alpha$ and $\beta$ on two datasets.
Sensitivity analysis of $m$.
To investigate the effect of latent space dimension, we conducted experiments by varying the parameter $m$ in the range of $[5,10,20,30,50]$. From Figure LABEL:m, we can observe that MemEvo exhibits significant robustness to dimensional changes and maintains stable performance across a wide range.
(b) The impact of $m$ and $\lambda$ on GRAZ02 dataset.
Sensitivity analysis of forgetting rate $\lambda$.
To investigate the impact of the forgetting rate $\lambda$, we conducted experiments on the GRAZ02 dataset, comparing different forgetting rate parameters within the range [0, 1, 1.5, 2]. As shown in Figure LABEL:lambda, when $\lambda=0$, the model degenerates to historical representation averaging, yielding the worst performance due to excessive smoothing of historical representation. Better performance occurs when $\lambda\in[1,2]$, indicating that controlled forgetting effectively filters redundant historical information while preserving useful knowledge.
Running time analysis.
To evaluate the computational efficiency of our method, we compare MemEvo with other incremental methods in Table 2 in terms of runtime on four datasets. Although LAIMVC performs excellently in terms of speed, MemEvo achieves a balance between performance and efficiency in all scenarios. Notably, unlike CMVC and FCMVC, which either fail to complete (OM) or require significantly more time on large-scale datasets, our method successfully handles challenging cases such as Caltech101-all and MNIST.
Times(s) GRAZ02 handwritten Caltech101-all MNIST
CMVC 0.44 0.63 35.56 53.36
LAIMVC 0.07 0.01 0.28 1.79
FDMVC 0.93 0.78 4.67 12.05
FCMVC 1.55 1.24 86.10 OM
Ours 0.22 0.17 2.00 18.45
Table 2: Running time comparison on four datasets.
Representation Visualization Analysis
To show the quality of the representations learned, we visualized the GRAZ02 dataset in Figure 5(a). As can be seen from the figure, our method performs better in terms of intra-class cohesion and inter-class dispersion. This is because our collaborative strategy allows the model to retain the historical core knowledge while learning new information, which is crucial for generating high-quality representations.
(a) T-SNE visualizations of learned representation.
Ablation experiments
To evaluate the contribution of each module to model performance, we conducted ablation experiments on four datasets in Table 3. Our experiments reveal the contribution of each module: using only reconstruction loss yields poor performance, confirming its limited discriminative power. Adding VAM brings modest gains, demonstrating its role in view alignment. KCM drives major improvements, proving its long-term knowledge consolidation capability. Furthermore, CFM also contributes to improvement, suggesting that selectively discarding outdated knowledge can benefit incremental tasks. The collaborative effect of the three modules enables the model to achieve better clustering performance.
Methods Flower17 GRAZ02 Caltech101-all MNIST
$\mathcal{L}_{recon}$ 40.00 38.25 17.06 35.78
$\mathcal{L}_{recon}$+VAM 42.13 38.75 17.29 37.74
$\mathcal{L}_{recon}$+KCM 87.06 77.37 62.03 60.55
Ours (w/o CFM) 91.47 85.50 63.81 88.10
Ours 94.85 92.28 67.10 99.87
Table 3: Ablation experiment results (w/o means without).
Impact of view increments
Figure 7(a) shows the trend in model performance as the number of views increases, with the single-view training results used as a baseline for comparison. Theoretically, as new views are introduced, clustering performance should exhibit a monotonically increasing trend. Experimental results indicate that MemEvo consistently shows gradual improvement as the number of views increases. Notably, our incremental method shows a significant advantage over single-view learning. This phenomenon indicates that our model not only effectively transfers knowledge from existing views but also continuously integrates information from new views.
(a) Impact of view number on two datasets.
Conclusion
This paper proposes an incremental multi-view clustering framework based on memory evolution. Our core contribution lies in designing a brain science-inspired component architecture that integrates cognitive forgetting, view alignment, and knowledge consolidation memory modules to achieve a balance between model plasticity and stability. Compared with current state-of-the-art incremental clustering methods, our model demonstrates significant advantages across multiple evaluation metrics, validating its strong potential for dynamic multi-view clustering tasks. Although the current model performs well in existing incremental scenarios, our model has not explored more challenging dynamic learning scenarios, such as class-incremental learning. In future research, we will develop an automatic class discovery model based on uncertainty estimation to eliminate dependence on the number of predefined classes.